OpenStudy (anonymous):

I need to define this explicitly: ln(3-x) - 1/2(ln(y^2+16)) = C Some one give me a hand just to make sure I did my algebra correctly

OpenStudy (amistre64):

for x or y ?

OpenStudy (anonymous):

For y

OpenStudy (amistre64):

\[\frac{ln(3-x) - 1}{2(ln(y^2+16))} = C\] this?

OpenStudy (anonymous):

Sorry lemme redo the equation. Lol it's a lot simpler than that

OpenStudy (anonymous):

\[\ln (3-x) - 1/2(\ln(y ^{2}+16)) = c \]

OpenStudy (amistre64):

\[ln (3-x) - ln(y ^{2}+16)^{1/2} = C\] \[ln (3-x) = C+ ln(y ^{2}+16)^{1/2}\] \[e^{ln (3-x)} = e^{C+ ln(y ^{2}+16)^{1/2}}\] \[3-x = e^C e^{ln(y ^{2}+16)^{1/2}}\] \[3-x = e^C (y^{2}+16)^{1/2}\] so far?

OpenStudy (anonymous):

Yeah looks good

OpenStudy (anonymous):

But wait, how did you go from C + ln(whatever) to multiplying them together?

OpenStudy (amistre64):

\[\left(3-x = e^C (y^{2}+16)^{1/2}\right)^2\] \[(3-x)^2 = e^{2C} (y^{2}+16)\] \[\frac{(3-x)^2}{e^{2C}} = y^2+16\] \[\frac{(3-x)^2}{e^{2C}}-16 = y^2\] \[\sqrt{\frac{(3-x)^2}{e^{2C}}-16} = y\]

OpenStudy (anonymous):

shouldn't it be e^C +(y^2+16)^1/2

OpenStudy (amistre64):

\[B^{a+b} = B^{a} B^{b}\]

OpenStudy (amistre64):

B=e a = C b = ln(...)

OpenStudy (anonymous):

so e^C + e^ln(y^2+16)^(1/2) = e^(C+(y^2+16)^(1/2)) ? That just doesn't quite seem right to me

OpenStudy (amistre64):

\[\sqrt{\frac{(3-x)^2}{e^{2C}}-16} = y\] \[\sqrt{\left(\frac{(3-x)}{e^{C}}\right)^2-(4)^2} = y\] can we go further?

OpenStudy (amistre64):

\[\huge e^{ln (3-x)} = e^{(C+ ln(y ^{2}+16)^{1/2})}\] \[\huge e^{ln (3-x)} = e^{C}\ e^{ln(y ^{2}+16)^{1/2}}\]

OpenStudy (anonymous):

Well C = ln(18) if what I did earlier in the problem was correct

OpenStudy (amistre64):

that does help :)

OpenStudy (amistre64):

when we "e" both sides; the RHS turns into a sum of exponents

OpenStudy (anonymous):

I have kind of a side question. does 1/2(ln(y^2+16)) = ln(y+4) ?

OpenStudy (amistre64):

im prone to say no; the sqrt of an sum of squares tends to be complex

OpenStudy (amistre64):

sqrt doesnt distribute over addition; only multiplication

OpenStudy (anonymous):

Well that's basically why I was getting wrong answers then because I put that in my problem haha

OpenStudy (amistre64):

\[\sqrt{ab}=\sqrt{a}*\sqrt{b}\] \[\sqrt{a+b}\ne\sqrt{a}+\sqrt{b}\]

OpenStudy (anonymous):

Gotcha, I shoulda known that

OpenStudy (anonymous):

I have a question, before you e'd both sides... wouldn't it have been easier to keep both logs on the same side? so you'd have like 3-x -1/2(y^2+16)?

OpenStudy (amistre64):

perhaps, but i never really go for "easier" :) it would have amounted to the same thing in the end yes